The function f: R^2~{(x,y)|y=0 or y= 5x} →R

defined by

f(x,y)= x^2-3xy/(y^2+5yx)

is a homogeneous function

True or false with full explanation


1
Expert's answer
2021-06-28T03:48:19-0400



TRUE



Explanation:



f(x, y)= "\dfrac{x^2 - 3xy }{y^2 + 5xy}"



To check whether f(x, y) is homogenous or not we replace x "\rightarrow" "\lambda"x and y "\rightarrow" "\lambda"y


and check if f( "\lambda"x, "\lambda"y) = f(x, y).




On replacing x "\rightarrow" "\lambda"x and y "\rightarrow" "\lambda"y, we have



f("\lambda"x, "\lambda"y) = "\dfrac{\lambda ^2x^2 - 3\lambda ^2xy }{\lambda ^2y^2 + 5\lambda ^2xy}"


Taking "\lambda ^2" common from numerator and denominator we have


f("\lambda"x, "\lambda"y) = "\dfrac{x^2 - 3xy}{y^2 + 5xy}" = f(x, y)


Hence, we see that f( "\lambda"x, "\lambda"y) = f(x, y). So, f(x, y) is homogenous.